PlEASE help! A postal clerk sold some fifteen-cent stamps and some twenty-five cent stamps. Altogether, 10 stamps were sold for a total cost of $1.70. How many of each type of stamp were sold?
Answer by Paul(988)

Let the $0.15 stamps be x
Let the $0.20 stamps be y
x+y=10
y=10-x

SInce I do not like working with demicals I'll convert them to integers by multiplying the numbers by 100.
So:
15x+20y=170 (subsituite the known value)
15x+20(10-x)=170
15x+200-20x=170
-5x=-30
x=6

y=10-6
y=4
Hence for $0.15 about 6 stamps were sold and for $0.20 about 4 stamps were sold.